A father is currently six times older than his son. In 10 years time his son will be 20 years younger than his dad. How old is the son now?
Answer by richwmiller(17219)

f=6s
s+10=f+10-20
s+10=f-10
s+20=f
s+20=6s
20=5s
s=4
f=24
check
14=34-20
ok